# ClaraVid: A Holistic Scene Reconstruction Benchmark From Aerial Perspective With Delentropy-Based Complexity Profiling

### v1

@@ 7 Delentropic Scene Profile Real Data
Figure 7: Reconstruction quality vs. scene complexity. The relationship between scene complexity, measured by the $\mu$ mean delentropy, and reconstruction fidelity (PSNR) across different neural reconstruction methods. A negative correlation is observed, where higher delentropy values correspond to lower PSNR scores, indicating increased reconstruction difficulty.
@@ 8.1 Reconstruction Performance
Throughout our experiments, we utilize the Nerfstudio framework[57] to systematically benchmark various neural reconstruction methods. For Nerfacto[57], we adopt the big configuration. TensoRF[56] is configured with an increased grid resolution of $500^{3}$. InstantNGP[36] employs 16 grid levels, a maximum resolution of 8192, and a hash map size of $2^{21}$. Zip-NeRF[58] is used in its default configuration without modifications. Gaussian Splatting[7] is initialized from the dataset’s scene-level point cloud with a spatial resolution of 100 cm, restricting point splitting beyond 3 million points to balance memory efficiency and reconstruction accuracy. Training durations vary across models, ranging from 60 to 720 minutes on an NVIDIA A6000 GPU, depending on architectural complexity. Across all experiments, we leverage FP16 training to optimize memory usage. The reported reconstruction and segmentation results correspond to a half-resolution output relative to the original image dimensions, maintaining a trade-off between computational feasibility and fidelity in large-scale aerial scene reconstruction.
@@ 8.2 Performance Across Varying Viewpoints
Both Nerfacto and Gaussian Splatting use the configuration from Section 8.2. For semantic learning, Nerfacto incorporates a jointly trained segmentation field, while Gaussian Splatting learns a separate set of 32-width semantic parameters—trained jointly but independent of geometry or positional encodings. Additionally, a 2-layer CNN classifier refines segmentation on the Gaussian splats. We compute per scene class weighting to balance semantic segmentation classes.
@@
9 Semantic Segmentation
We train a DPT[61] model with a DinoV2-L[62] backbone on both synthetic and real-world UAV datasets. The backbone is kept frozen while only the decoder is optimized. Our training protocol is standardized across datasets, with the number of steps adjusted proportionally to the dataset size—ranging from 5,000 to 30,000—to prevent overfitting. We use each dataset’s native resolution and perform random crops of 630×630 pixels. Additionally, we uniformly apply data augmentations such as color adjustments, lighting variations, rotations, and distortions. Finally, the semantic palette is adapted to the real dataset following the [12] protocol.
10 Claravid
In this section, we present additional insights into Claravid.
@@ 10.2 Rendering Pipeline Configuration
To achieve high-fidelity rendering of large-scale aerial scenes in Unreal Engine 4 (UE4), we implement targeted modifications to the rendering pipeline, prioritizing fine detail preservation and visibility consistency at extended distances. To mitigate the disappearance of distant objects and enhance detail clarity, we increase the rendering resolution by setting r.ScreenPercentage=300, effectively supersampling the scene at 3 times the native resolution. This adjustment minimizes aliasing artifacts—particularly pronounced in oblique aerial perspectives—and ensures that small-scale features remain discernible, albeit at a higher computational cost justified by the resulting visual fidelity. Shadow integrity for distant fine elements, such as foliage and thin structures, is preserved by reducing r.Shadow.MinRadius to 0.001, which enables shadow casting for geometrically narrow features that would otherwise be lost. Shadow map resolution is also increased to 2048×2048 for objects considered having fine details, maintaining sharp shadow edges across expansive terrains. Geometric fidelity is ensured by disabling level-of-detail (LOD) transitions, preventing mesh simplification that degrades structural complexity at greater distances or altitudes, a choice prioritizing rendering quality over real-time performance scalability. Similarly, we disable distance-based object culling to maintain persistent visibility of all scene elements, eliminating sudden visibility discontinuities that disrupt spatial and temporal coherence critical for synthetic aerial datasets. We present a visual overview of the dataset in Section 10.
Figure 10: Claravid Overview. Representative images across 5 different scenes that were either created or improved. From top to bottom: Urban Dense, Nature, Rural, Urban, Highway.

### v2

@@ 9.2 DSP Analysis on Large Reconstructions
Figure 9: Reconstruction quality vs. scene complexity. The relationship between scene complexity, measured by the $\mu$ mean delentropy, and reconstruction fidelity (PSNR) across different neural reconstruction methods. A higher delentropy value correspond to lower PSNR scores, indicating increased reconstruction difficulty.
@@ 10.1 Reconstruction Performance
Throughout our experiments, we utilize the Nerfstudio framework[62] to systematically benchmark various neural reconstruction methods. For Nerfacto[62], we adopt the big configuration. TensoRF[61] is configured with an increased grid resolution of $500^{3}$. InstantNGP[40] employs 16 grid levels, a maximum resolution of 8192, and a hash map size of $2^{21}$. Zip-NeRF[63] is re-implemented and used with default configuration. Gaussian Splatting[7] is initialized from the dataset’s scene-level point cloud with a spatial resolution of 100 cm, restricting point splitting beyond 3 million points to balance memory efficiency and reconstruction accuracy. Training durations vary across models, ranging from 60 to 720 minutes on an NVIDIA A6000 GPU, depending on architectural complexity. Across all experiments, we leverage FP16 training to optimize memory usage. The reported reconstruction and segmentation results correspond to a half-resolution output relative to the original image dimensions, maintaining a trade-off between computational feasibility and fidelity in large-scale aerial scene reconstruction.
@@ 10.2 Performance Across Varying Viewpoints
Both Nerfacto and Gaussian Splatting use the configuration described previously in Supplementary 10.1. For semantic learning, Nerfacto incorporates a jointly trained segmentation field, while Gaussian Splatting learns a separate set of 32-width semantic parameters—trained jointly but independent of geometry or positional encodings. Additionally, a 2-layer CNN classifier refines segmentation on the Gaussian splats. We compute per scene class weighting to balance semantic segmentation classes. The depth for both methods is obtained from rendering with no additional heads or learned parameters.
10.3 Semantic Segmentation
We train a DPT[69] model with a DinoV2-L[70] backbone on both synthetic and real-world UAV datasets. The backbone is kept frozen while only the decoder is optimized. Our training protocol is standardized across datasets, with the number of steps adjusted proportionally to the dataset size—ranging from 5,000 to 30,000—to prevent overfitting. We use each dataset’s native resolution and perform random crops of 630×630 pixels. Additionally, we uniformly apply data augmentations such as color adjustments, lighting variations, rotations, and distortions. Finally, the semantic palette is adapted to the real dataset following the [12] protocol, where both training and evaluation are done using reduced pallette.
10.4 UAVid Evaluation
For the UAVid evaluation, we focus on mostly static scenes, as dynamic object masks are not available. We select six representative sequences (13, 15, 29, 31, 36, 38), each covering an area of approximately $\approx$ 0.1 km2. All sequences consist of 901 frames, training samples are selected at every fourth index, while testing ones are chosen at every second offset (i.e., $i\equiv 2\pmod{4}$). Stationary frames with minimal motion are excluded, as they often are poorly registered. Additionally, we consider the first and last 100 frames in train set, to mitigate any potential collection policy errors and minimize the spatial content outside the region of interest. All images are downsampled by a factor of 2, and reconstructions are performed using COLMAP[71, 72]. All other settings for the evaluated models follow the protocol described in Section 10.1.
11 Claravid
In this section, we present additional insights into Claravid. Additionally we present a more detailed visual overview in Figure 12.
@@ 11.1 Semantic Complexity
Figure 12: ClaraVid Modalities. Overview of the 8 UAV missions, showing representative frames with RGB, depth, semantic, instance, and dynamic annotations, along with scene-level point cloud views.
@@ 11.2 Rendering Pipeline Configuration
To achieve high-fidelity rendering of large-scale aerial scenes in Unreal Engine 4, we implement targeted modifications to the rendering pipeline, prioritizing fine detail preservation and visibility consistency at extended distances. To mitigate the disappearance of distant objects and enhance detail clarity, we increase the rendering resolution by setting r.ScreenPercentage=300, effectively supersampling the scene at 3 times the native resolution. This adjustment minimizes aliasing artifacts—particularly pronounced in oblique aerial perspectives—and ensures that small-scale features remain discernible, albeit at a higher computational cost justified by the resulting visual fidelity. Shadow integrity for distant fine elements, such as foliage and thin structures, is preserved by reducing r.Shadow.MinRadius to 0.001, which enables shadow casting for geometrically narrow features that would otherwise be lost. Shadow map resolution is also increased to 2048×2048 for objects considered having fine details, maintaining sharp shadow edges across expansive terrains. Geometric fidelity is ensured by disabling level-of-detail (LOD) transitions, preventing mesh simplification that degrades structural complexity at greater distances or altitudes, a choice prioritizing rendering quality over real-time performance scalability. Similarly, we disable distance-based object culling to maintain persistent visibility of all scene elements, eliminating sudden visibility discontinuities that disrupt spatial and temporal coherence.